### v1

@@ 3.2 Determination of instrumental resolution and threshold
We measured the instrumental resolution ($\sigma_{\text{inst}}$) or LSF for each galaxy using its respective variance spectrum. In this regard, we fitted Gaussian profiles plus a constant (for background noise) to the isolated and unblended sky lines at different wavelengths. The fitted second moment of a Gaussian profile is considered $\sigma_{\text{inst}}$ at that wavelength. The $\sigma_{\text{inst}}$ is first measured in Å and then converted to km s-1 using $\sigma_{\text{inst}}=\frac{\Delta\lambda c}{\lambda}$. In Figure 5, we show how $\sigma_{\text{inst}}$ varies typically as a function of wavelength for three instruments – LRIS-red, Xshooter, and DEIMOS. As $\sigma_{\text{inst}}$ for all three instruments does not change significantly over the fitted wavelength range, we use the respective average $\sigma_{\text{inst}}$ within the fitted wavelength range of each galaxy for kinematics measurements. We report the average $\sigma_{\text{inst}}$ of each galaxy in Table LABEL:table:veldis_measurements. The associated uncertainties on $\sigma_{\text{inst}}$ are the sample standard deviations of the measured $\sigma_{\text{inst}}$ at different wavelengths.
@@
We impose a criterion to ensure that any bias due to imprecise measurement of $\sigma_{\text{inst}}$ is less than 1%. Following KM25 this requires:
@@
Thus, depending on the measured $\sigma$, we can only accept instrumental setups for which the criterion is satisfied. We found that all 47 galaxies listed in Table LABEL:table:veldis_measurements satisfy this requirement.
3.3 Determination of SNR threshold
@@
It is well known that spectra with insufficient SNR could result in biased kinematics (e.g., Treu et al. 2001). To determine the SNR required for the unbiased measurements below 1%, we performed numerical simulations with each specific instrumental setup (resolution, sampling, etc.) used to collect the galaxy spectra of the SL2S sample. We conducted simulations for five setups: LRIS-blue, LRIS-red low-resolution, LRIS-red high-resolution, DEIMOS, and Xshooter. We started by creating mock galaxy spectra using G/K-type stars from the Indo-US and MILES stellar library. We resampled the selected template spectra linearly as per the reciprocal dispersion of a particular instrument setup and convolved the spectra with a Gaussian kernel using pPXF. The $\sigma$ value of the Gaussian kernel is determined by the velocity dispersion to be injected and the instrumental resolution, $\sigma_{\text{inst}}$, of the setup. Then we added noise to the convolved spectra to mimic a particular SNR. The reciprocal dispersions we used for these setups are 1.45, 1.85, 0.8, 0.65, 0.2 Å respectively. Similarly, the instrumental resolutions applied on the mock galaxy spectra for the setups are $\sigma_{\text{inst}}=$ 200, 200, 80, 70, 35 km s-1 respectively. The chosen reciprocal dispersions and $\sigma_{\text{inst}}$ are representative of the actual SL2S data collected by each setup. For a range of velocity dispersions values ( $\sigma_{\text{input}}$ = 150, 200, 250, 300 km s-1) and SNR (10, 15, 20, 25, 30, 35, 40), we created 1000 mock galaxy spectra and recovered the input velocity dispersion using clean Indo-Us stellar library and the pPXF package [^3] https://pypi.org/project/ppxf/ (Cappellari 2017, 2023).
@@
From these simulations, we found that, for LRIS-red high-resolution, DEIMOS, and Xshooter, a SNR threshold of 15 per Å is sufficient to obtain an unbiased estimate below 1%. However, LRIS-blue and LRIS-red low-resolution setups do not have sufficient spectral resolution to guarantee bias below 1% even at high SNR and velocity dispersion. Additionally, as the $\sigma_{\text{inst}}$ of these two setups are comparable to the velocity dispersions of the galaxies to be measured ($\sim$ 200 km s-1), we decided to exclude the spectra from these two setups in this paper. Fortunately, we have to discard only two systems from the sample due to this constraint. In all other cases there are complementary data available from the LRIS-red high-resolution setup.
3.4 Measuring SNR of the spectra
@@
We measured the SNR per Å of all available spectra from the three accepted instrument setups described in the last section. We used three different rest frame wavelength intervals — 4000–4400, 4800–5200, and 5400–5800 Å, to determine the average SNR per pixel of the spectra and divided it by the square root of the number of Å per pixel to obtain the average SNR per Å. We selected these three wavelength regions as they have relatively uniform SNR per pixel. If a spectrum does not contain all three regions, then we use whatever wavelength ranges from these three regions are available. We excluded pixels covering atmospheric A and B lines if present in these three regions while measuring average SNR. We found 30 galaxies with spectra having SNR per Å $\geq$ 15. However, we still measured and reported kinematics and associated uncertainties of the other galaxies as they could be used in projects where bias above 1% is acceptable. The SNR of the fitted galaxy spectra is listed in Table LABEL:table:veldis_measurements.
@@ 3.5 Selection of clean stellar libraries
KM25 discussed extensively the importance of using clean stellar templates and showed that the usage of defective templates can introduce bias in the kinematics measurements. Therefore, we have created two clean stellar libraries– Indo-Us and MILES using the list of clean stars provided in KM25. Some of the stars in these lists also contain defects in several wavelength regions, and KM25 further flagged these stars with the information of the problematic wavelength range. We checked for the presence of these defective regions in the clean templates within the wavelength range of 3600-6000 Å. If defects are present, we also excluded these spectra from the clean library. We used the clean Indo-US and MILES library for all the kinematics measurements and systematic tests. The final clean Indo-US and MILES libraries contain 802 and 789 templates, respectively.
We initially also considered using the Xshooter stellar library. However, more than 50% of the clean star templates in this library have been flagged beyond 5300 Å due to artifacts around the dichroic separating the UVB from the visible wavelength range, while for most galaxies we have to fit up to 6000 Å at rest-frame. As the clean Xshooter library has to be downsized significantly beyond 5300 Å, we decided not to use this library in this project.
@@ 3.6 Determination of polynomial orders
Following the recipe in KM25, we ran a grid of polynomial orders to determine the optimal orders of the additive and multiplicative polynomials that would make the kinematics measurements stable and reduce the systematic scatter in measurements using different template libraries for the whole sample. A similar test was conducted by KM25 using a sample of galaxy spectra covering the same rest-frame wavelength range and observed with the same instrument. However, the galaxy spectra in the SL2S sample contain different rest-frame wavelength ranges and were collected using different instruments. Therefore, we also checked whether the optimal polynomial orders depend on the instrument or the fitted wavelength range. We found that higher polynomial orders, such as 7-9 for additive polynomials and 2 for multiplicative polynomials, reduce the systematic scatter regardless of the instrument or wavelength range. We decided to use orders of 8 and 2, for additive and multiplicative polynomials, respectively. As a test, we also computed what would be the additional scatter in stellar velocity dispersion if we marginalized over polynomial orders around the minimal choice, and we found it to be much less than 0.5%, therefore negligible with respect to that attributable to stellar templates.
@@ 3.7 Wavelength range and masking
The galaxy spectra of the SL2S sample cover different rest-frame wavelength ranges due to combination of redshift and instrument setup. Regardless, we used the maximum possible wavelength range available for a galaxy spectrum without compromising the SNR. The cutoff of the adopted wavelength range on the blue side is imposed both by the instrument setup and stellar template coverage, and on the red side by the redshifts of the galaxies. In Figures 1 - 4, we show the fitted wavelength range for each galaxy. We excluded some wavelength regions from the fits at two stages. First, we masked the rest-frame wavelength range 6850-6950 Å and 7580-7750 Å  corresponding to the atmospheric B-band and A-band, respectively, if present. Next, using an iterative sigma clipping method, we detected wavelength ranges containing emission lines, instrument or extraction related artifacts, cosmic rays, etc., if any, and also masked them from the fit. In Figures 1 - 4, these masked regions are shown in gray bands.
[…]
@@
As an additional test, we repeated the measurement for a subset of galaxies with spectra with SNR $>20$/Å, masking at random one of the most prominent absorption features present in the spectra (CaHK, Mgb, G-band, NaD). The results are shown in Figure 6, separately for Indo-US and MILES library, in order to isolate the effect of the lines from that of the stellar library. On average the stellar velocity dispersion changes by less than 1%, demonstrating that for data and templates of sufficient quality the results do not depend on any single feature.
@@
Figure 6: For each galaxy with high-quality spectra in the sample (SNR$>20$/Å) we show the mean and scatter in the value of stellar velocity dispersion when masking one of the main absorption features (CaHK, G-band, MgB, NaD) with respect to the value obtained for the entire spectrum. Thick horizontal lines represent the sample average and dashed lines the error on the mean.
3.8 Determining the threshold to count effective template number
For the reasons discussed earlier, we used two clean stellar libraries for kinematic fitting and combined the measurements and associated uncertainties as a weighted average. The weight assigned to each template library is determined using BIC for the whole sample. The total BIC and associated weight is calculated following equations 5-10 in KM25. One crucial step in measuring BIC is to determine the effective number of templates that provide a significant contribution to the kinematic model. KM25 extensively discussed the issue and how to address it. Briefly, in an example of overfitting the noise, the model could include some templates with very low weights which may improve the fit (lowering $\chi^{2}$) by a minimal amount. However, they degrade the BIC significantly by increasing the total number of parameters. This effect can lead to incorrect estimates of the weights for the template libraries. Following the procedure in KM25, we first carried out a pPXF fit with all the templates in a clean library, and then performed successive fits while only using the subset of templates that contribute only above a certain threshold to the initial fit. The threshold is the weight of the templates in defining the model for the initial fit. As expected, we find that for a higher threshold (10-20%), the BIC is relatively large because there are not enough templates to obtain a good fit. However, below $\sim 5$%, the $\chi^{2}$ does not improve enough to compensate for the increased number of degrees of freedom in the BIC. Therefore, we adopted a 5% threshold to count the effective number of templates for the SL2S sample (as expected the threshold is similar to the inverse of the SNR of the dataset, see KM25).
@@ 3.8 Determining the threshold to count effective template number
Figure 7: BIC-weighted velocity dispersions of the SL2S sample as a function of redshift. The green, red, and black markers show galaxies in three different SNR bins - SNR ¿ 20, 13 ¡ SNR $\leq$ 20, and SNR ¿20, respectively. The associated error bars show total uncertainties in the measured velocity dispersions and are dominated by the statistical uncertainties, which are driven by the SNRs.
@@ 4 New measurements and uncertainties
In this section, we present the results obtained by using the methodology discussed in the previous section. The fitted models from pPXF are shown in Figures 1 to 4, and new measurements and uncertainties are summarized in Table LABEL:table:veldis_measurements. For each galaxy, we report the SNR and instrumental resolution of the spectrum, our best estimate of the stellar velocity dispersion, together with the random and systematic errors. The mean $\bar{\sigma}$, the systematic uncertainty $\Delta\bar{\sigma}$, and statistical uncertainty $\delta\bar{\sigma}$ of the velocity dispersion of an individual galaxy are defined as in KM25 such that
@@
where the index $k$ runs through the template library, $\delta\sigma_{k}$ is the statistical uncertainty estimated by pPXF, and $w_{k}$ is the weight assigned.
Following KM25, we provide the stellar velocity dispersions and uncertainties as both the BIC-weighted average and equally-weighted average of the measurements from two template libraries. The BIC determined weights were estimated at the sample level to minimize noise in the process, and the samples were grouped by instrument and/or SNR. The estimated weights for the groups are summarized in Table 2. Note that preference for one of the libraries changes depending on the instrument and the SNR. For each galaxy, we used the weight for the group it is affiliated with to measure BIC-weighted average values. For example, if a galaxy is observed with LRIS and has an SNR$>$20/Å, then the assigned weights to the measurements from Indo-US and MILES are 0.78 and 0.22, respectively. In general, the BIC-weighted and equally-weighted velocity dispersions are very similar, consistent with the good agreement between the template libraries.
When the number of template libraries is small (as in the case considered here), it is prudent to adopt Bessel’s correction to get the unbiased estimators (see KM25 for a derivation). Thus, the systematic uncertainties are Bessel corrected following:
@@
However, for $k=2$, the Bessel correction nullifies the dependence on the weights and $\Delta_{B}\bar{\sigma}$ reduces to $(\Delta_{B}\bar{\sigma})^{2}=(\frac{\sigma_{1}-\sigma_{2}}{2})^{2}$ where $\sigma_{1}$, and $\sigma_{2}$ are measurements using two template libraries. Thus, the Bessel-corrected systematic uncertainties, $\Delta_{B}\bar{\sigma}$, for the BIC-weighted and equally-weighted velocity dispersions are equal and reported as a single column in Table LABEL:table:veldis_measurements.
We have also measured the BIC weighted average statistical and systematic uncertainty, both Bessel corrected and non-corrected, for subsamples grouped by SNR and/or instrumental setup. In addition, we estimate the covariance matrix as follows:
@@
where the indices $i,j$ identify the elements of the covariance matrix, $\bar{\sigma}_{i,j}$ is the weighted average velocity dispersion of the element at index $i$ or $j$, and $\delta_{i,j}$ is the Kronecker delta function. And as above, the index $k$ runs through the template library, $\delta\bar{\sigma}_{i}$ is the weighted average statistical uncertainty, and $w_{k}$ is the weight assigned. Along the diagonal, the first term is the systematic error, and the second is the statistical error. With Bessel correction, the estimator reads as
@@
The normalized sample average uncertainties and the average amplitude of the off-diagonal terms of the covariance matrix between elements are summarized in Table 2. As expected, the random error is driven by the SNR, regardless of the instrumental setup, and ranges between 3-8%. The systematic errors also depend on the SNR and instrumental setup, and are at the level of 1-2% without Bessel correction and around 2-3% with the correction. This is, of course, only the error associated with template choice. Other potential sources of error, such as the wavelength range and polynomial order, were ignored, as they are negligible with respect to the template contribution.
The average off-diagonal amplitude of the covariance matrix for different subsamples is positive. This is expected as the templates yielding higher (lower) velocity dispersion do so consistently for all galaxies. This quantity ranges around 1-2% within the same instrumental setup, and subpercent across different instrumental setups. The LRIS subsamples have slightly higher covariances than the combined Xshooter and DEIMOS samples.
We found that Indo-US library is preferred by the LRIS data, while MILES is preferred by the other. This emphasizes the importance of using several stellar libraries and calculating BIC-weighted averages. However, in this case, similar results are obtained if the templates are given equal weight, so the difference in BIC weight has less impact.
As a consistency test of our error estimates, we compare the stellar velocity dispersion obtained for 11 systems that have multiple spectra with sufficient resolution. The fiducial spectrum is the one with the highest SNR. The other spectra may have significantly lower SNR and may cover wavelength ranges differing by 500-1000 Å from their fiducial counterpart. The difference in the SNR quality of the spectra is reflected in the random uncertainties listed in Table LABEL:table:veldis_measurements. The comparison is shown in Figure 8, with the error bars representing the sum in quadrature of the random and systematic errors. The sample average of the ratio is $<\frac{\sigma_{\mathrm{high\ SNR}}}{\sigma_{\mathrm{low\ SNR}}}-1>=-0.11\pm 2.70$. Repeated measurements with higher SNRs are needed to carry out an even more stringent test of systematics across different instrumental setups.

### v2

@@ 3.2 Determining the instrumental resolution and threshold
We measured the instrumental resolution ($\sigma_{\text{inst}}$) or LSF for each galaxy using its respective variance spectrum. To do this, we fit Gaussian profiles plus a constant (for the background noise) to the isolated and unblended sky lines at different wavelengths. The second moment of the fitted Gaussian profile was considered $\sigma_{\text{inst}}$ at that wavelength. The $\sigma_{\text{inst}}$ was first measured in Å and then converted into km s-1 using $\sigma_{\text{inst}}=\frac{\Delta\lambda c}{\lambda}$. In Fig. 5 we show the typical wavelength dependence of $\sigma_{\text{inst}}$ for three instruments: LRIS-red, Xshooter, and DEIMOS. As $\sigma_{\text{inst}}$ for all three instruments does not change significantly over the wavelength range we fit, we used the respective average $\sigma_{\text{inst}}$ within the fitted wavelength range of each galaxy for kinematics measurements. We report the average $\sigma_{\text{inst}}$ of each galaxy in Table LABEL:table:veldis_measurements. The associated uncertainties on $\sigma_{\text{inst}}$ are the sample standard deviations of the measured $\sigma_{\text{inst}}$ at different wavelengths.
@@
We imposed a criterion to ensure that any bias due to imprecise measurement of $\sigma_{\text{inst}}$ was lower than 1%. Following KM25, this required
@@
Thus, depending on the measured $\sigma$, we only accepted instrumental setups for which the criterion was satisfied. We found that all 47 galaxies listed in Table LABEL:table:veldis_measurements satisfied this requirement.
3.3 Determining the S/N threshold
@@
It is well known that spectra with an insufficient S/N can result in biased kinematics (e.g., Treu01). To determine the S/N required for the unbiased measurements below 1%, we performed numerical simulations with each specific instrumental setup (resolution, sampling, and so on) used to collect the galaxy spectra of the SL2S sample. We conducted simulations for five setups: LRIS-blue, LRIS-red low-resolution, LRIS-red high-resolution, DEIMOS, and Xshooter. We started by creating mock galaxy spectra using G/K-type stars from the Indo-US and MILES stellar library. We resampled the selected template spectra linearly as per the reciprocal dispersion of a particular instrument setup and convolved the spectra with a Gaussian kernel using pPXF. The $\sigma$ value of the Gaussian kernel was determined by the velocity dispersion to be injected and the instrumental resolution, $\sigma_{\text{inst}}$, of the setup. Then, we added noise to the convolved spectra to mimic a particular S/N. The reciprocal dispersions we used for these setups were 1.45, 1.85, 0.8, 0.65, and 0.2 Å respectively. Similarly, the instrumental resolutions applied on the mock galaxy spectra for the setups were $\sigma_{\text{inst}}=$ 200, 200, 80, 70, and 35 km s-1 respectively. The chosen reciprocal dispersions and $\sigma_{\text{inst}}$ are representative of the actual SL2S data collected by each setup. For a range of velocity dispersion values ( $\sigma_{\text{input}}$ = 150, 200, 250, and 300 km s-1) and S/N (10, 15, 20, 25, 30, 35, and 40), we created 1000 mock galaxy spectra and recovered the input velocity dispersion using the clean Indo-Us stellar library and the pPXF package [^3] https://pypi.org/project/ppxf/ (Cappellari_2017_ppxf; LEGAC_ppxf_Cappellari_2023).
@@
Based on these simulations, we found that for LRIS-red high-resolution, DEIMOS, and Xshooter, a S/N threshold of 15 per Å was sufficient to obtain an unbiased estimate below 1%. The spectral resolution of the LRIS-blue and LRIS-red low-resolution setups is not sufficient to guarantee a bias below 1% even at a high S/N and velocity dispersion, however. Additionally, as the $\sigma_{\text{inst}}$ of these two setups are comparable to the velocity dispersions of the galaxies to be measured ($\sim$ 200 km s-1), we decided to exclude the spectra from these two setups in this paper. Fortunately, we only had to discard two systems from the sample because of this constraint. In all other cases, complementary data were available from the LRIS-red high-resolution setup.
3.4 Measuring the S/N of the spectra
@@
We measured the S/N per Å of all available spectra from the three accepted instrument setups described in the last section. We used three different rest frame wavelength intervals, 4000–4400, 4800–5200, and 5400–5800 Å, to determine the average S/N per pixel of the spectra and divided it by the square root of the number of Å per pixel to obtain the average S/N per Å. We selected these three wavelength regions as they have relatively uniform S/N per pixel. When a spectrum did not contain all three regions, we used whatever wavelength ranges from these three regions were available. We excluded pixels covering the atmospheric A and B lines if present in these three regions while measuring the average S/N. We found 30 galaxies with spectra that had an S/N per Å $\geq$ 15. We still measured and report the kinematics and associated uncertainties of the other galaxies, however, because they might be used in projects where a bias above 1% is acceptable. The S/N of the fitted galaxy spectra is listed in Table LABEL:table:veldis_measurements.
@@ 3.5 Selection of clean stellar libraries
KM25 extensively discussed the importance of using clean stellar templates and showed that the usage of defective templates can introduce bias in the kinematics measurements. Therefore, we created two clean stellar libraries, Indo-Us and MILES, using the list of clean stars provided by KM25. Some of the stars in these lists also contain defects in several wavelength regions, and KM25 further flagged these stars with the information of the problematic wavelength range. We checked for the presence of these defective regions in the clean templates within the wavelength range of 3600-6000 Å. When defects were present, we also excluded these spectra from the clean library. We used the clean Indo-US and MILES library for all the kinematics measurements and systematic tests. The final clean Indo-US and MILES libraries contained 802 and 789 templates, respectively.
We initially also considered using the Xshooter stellar library. More than 50% of the clean star templates in this library were flagged beyond 5300 Å because of artifacts around the dichroic that separates the UVB from the visible wavelength range, however, while for most galaxies, we had to fit up to 6000 Å at rest-frame. As the clean Xshooter library has to be downsized significantly beyond 5300 Å, we decided not to use this library in this project.
@@ 3.6 Determination of polynomial orders
Following the recipe in KM25, we ran a grid of polynomial orders to determine the optimal orders of the additive and multiplicative polynomials that would make the kinematics measurements stable and reduce the systematic scatter in measurements using different template libraries for the whole sample. A similar test was conducted by KM25 using a sample of galaxy spectra that covered the same rest-frame wavelength range and was observed with the same instrument. The galaxy spectra in the SL2S sample contain different rest-frame wavelength ranges and were collected using different instruments, however. Therefore, we also checked whether the optimal polynomial orders depend on the instrument or the fitted wavelength range. We found that higher polynomial orders, such as 7-9 for additive polynomials and 2 for multiplicative polynomials, reduce the systematic scatter regardless of the instrument or wavelength range. We decided to use orders of 8 and 2 for additive and multiplicative polynomials, respectively. As a test, we also computed the additional scatter in stellar velocity dispersion if we marginalized over polynomial orders around the minimal choice, and we found it to be much lower than 0.5%, which is negligible with respect to that attributable to stellar templates.
@@ 3.7 Wavelength range and masking
The galaxy spectra of the SL2S sample cover different rest-frame wavelength ranges due to combination of redshift and instrument setup. Regardless, we used the widest possible wavelength range available for a galaxy spectrum without compromising the S/N. The cutoff of the adopted wavelength range on the blue side was imposed by the instrument setup and the stellar template coverage, and on the red side by the redshifts of the galaxies. In Figs. 1 to4, we show the wavelength range we fit for each galaxy. We excluded some wavelength regions from the fits at two stages. First, we masked the rest-frame wavelength range 6850-6950 Å and 7580-7750 Å  corresponding to the atmospheric B and A band, respectively, if present. Next, using an iterative sigma clipping method, we detected wavelength ranges containing emission lines, instrument or extraction related artifacts, cosmic rays, and so on, if any, and also masked them from the fit. In Figs. 1 - 4, these masked regions are shown in gray bands.
Following KM25, we also tested for systematic bias due to small changes (100-200 Å) in the fitted wavelength range. For this test, we selected eight galaxies with the longest wavelength coverage from three instruments. Those with Xshooter and DEIMOS spectroscopy have rest-frame wavelength ranges of about 3600 - 6000 Å  and those from LRIS-red have a wavelength coverage of about 4400 - 6000 Å. This long wavelength range is conducive to making small changes on the fitted wavelength range on the blue and red sides while keeping the prominent absorption features such as CaHK, Mgb, or NaD, if present. We trimmed and/or extended the original wavelength range on either end by 100 - 200 Å and fit for kinematics using the same setting as for the fiducial wavelength range. We then calculated the average deviation in the measured velocity dispersions due to wavelength change relative to the respective fiducial velocity dispersion for each galaxy and for all of them together. We found that the sample average deviation is $\overline{\left<\frac{\sigma_{\lambda_{k}}}{\sigma_{0}}-1\right>_{i}}=0.8\pm 0.5$%. Here, $i$ represents each of the eight galaxy in the sample, $\sigma_{0}$ is the velocity dispersion from the fiducial wavelength range, and $\sigma_{\lambda_{k}}$ are the velocity dispersions from the modified wavelength ranges. As the uncertainty from the changes in wavelength is subdominant to the total error budget at the sample level, we ignored this in estimating the final systematic uncertainty.
@@
As an additional test, we repeated the measurement for a subset of galaxies with spectra with an S/N >20>20/Å, masking at random one of the most prominent absorption features in the spectra (CaHK, Mgb, G-band, or NaD). The results are shown in Fig. 6 separately for the Indo-US and MILES library, in order to isolate the effect of the lines from that of the stellar library. The stellar velocity dispersion changes by less than 1% on average, which demonstrates that the results for data and templates of sufficient quality do not depend on any single feature.
@@
Figure 6: For each galaxy with high-quality spectra in the sample (S/N>20>20/Å), we show the mean and scatter in the value of stellar velocity dispersion when masking one of the main absorption features (CaHK, G-band, MgB, or NaD) with respect to the value obtained for the entire spectrum. The thick horizontal lines represent the sample average, and the dashed lines show the error on the mean.
@@
Figure 7: BIC-weighted velocity dispersions of the SL2S sample as a function of redshift. The green, red, and black markers show galaxies in three different S/N bins, S/N ¿ 20, 13 ¡ S/N $\leq$ 20, and S/N ¿20, respectively. The associated error bars show the total uncertainties in the measured velocity dispersions and are dominated by the statistical uncertainties, which are driven by the S/Ns.
3.8 Determining the threshold to count the effective template number
For the reasons discussed above, we used two clean stellar libraries for the kinematic fitting and combined the measurements and associated uncertainties as a weighted average. The weight assigned to each template library was determined using the BIC for the whole sample. The total BIC and associated weight were calculated following Eqs. (5)-(10) in KM25. One crucial step in measuring BIC is to determine the effective number of templates that provide a significant contribution to the kinematic model. KM25 extensively discussed the issue and how to address it. Briefly, in an example of overfitting the noise, the model could include some templates with very low weights that might improve the fit (lowering $\chi^{2}$) by a minimal amount. They degrade the BIC significantly by increasing the total number of parameters, however. This effect can lead to incorrect estimates of the weights for the template libraries. Following the procedure in KM25, we first carried out a pPXF fit with all the templates in a clean library, and we then performed successive fits while only using the subset of templates that only contribute above a certain threshold to the initial fit. The threshold is the weight of the templates in defining the model for the initial fit. As expected, for a higher threshold (10-20%), the BIC is relatively large because there are not enough templates to obtain a good fit. Below $\sim 5$%, however, the $\chi^{2}$ does not improve enough to compensate for the increased number of degrees of freedom in the BIC. Therefore, we adopted a 5% threshold to count the effective number of templates for the SL2S sample (as expected, the threshold is similar to the inverse of the S/N of the dataset; see KM25).
Table 1: BIC-weighted uncertainties for the whole SL2S sample and subsamples. [^4] We report the BIC-weighted average statistical uncertainty, the average systematic uncertainty (both Bessel corrected and uncorrected), and the average amplitude of the off-diagonal terms of the covariance matrix between elements of the sample (both Bessel corrected and uncorrected) for the whole SL2S sample and subsamples grouped by instrument and S/N. The BIC-determined weights given to the two libraries for each subsample and the number of galaxies in each subsample are also provided.
Sample $\delta\bar{\sigma}/\bar{\sigma}> <$$\Delta\bar{\sigma}/\bar{\sigma}> <$$\sqrt{<C_{i,j}\bar{\sigma}/\bar{\sigma_{i}}\bar{\sigma_{j}}>_{i\neq j}}$ $\Delta_{B}\bar{\sigma}/\bar{\sigma}> <$$\sqrt{<C_{B,i,j}\bar{\sigma}/\bar{\sigma_{i}}\bar{\sigma_{j}}>_{i\neq j}}$ Indo-US MILES Num. of Obj.
SL2S S/N ¿ 20 3.47% 0.66% 0.22% 2.14% 0.72% 0.05 0.95 15
SL2S 13 ¡ S/N $\leq$ 20 4.60% 1.07% 0.65% 1.70% 1.04% 0.27 0.73 18
SL2S S/N $\leq$ 13 7.40% 1.68% 0.90% 3.07% 1.65% 0.81 0.19 14
SL2S All 5.05% 1.38% 0.67% 2.22% 1.08% 0.26 0.74 47
SL2S-LRIS S/N ¿ 20 3.73% 1.39% 1.07% 2.38% 1.83% 0.78 0.22 8
SL2S-LRIS 13 ¡ S/N $\leq$ 20 4.68% 1.42% 0.97% 2.09% 1.42% 0.37 0.63 13
SL2S-LRIS S/N $\leq$ 13 6.90% 0.91% 0.43% 2.73% 1.30% 0.94 0.06 10
SL2S-LRIS All 5.13% 1.63% 1.02% 2.36% 1.48% 0.62 0.38 31
SL2S-Other S/N ¿ 20 3.16% 0.003% 0.001% 1.97% 0.45% 0.001 0.999 7
SL2S-Other 13 ¡ S/N $\leq$ 20 4.38% 0.40% 0.04% 0.71% 0.06% 0.19 0.81 5
SL2S-Other S/N $\leq$ 13 8.68 % 2.49% 1.55% 3.69% 2.30% 0.35 0.65 4
SL2S-Other All 4.94% 0.10% 0.02% 1.96% 0.35% 0.001 0.999 16
@@ 4 New measurements and uncertainties
In this section, we present the results we obtained by using the method discussed in the previous section. The fitted models from pPXF are shown in Figs. 1 to 4, and the new measurements and uncertainties are summarized in Table LABEL:table:veldis_measurements. For each galaxy, we report the S/N and instrumental resolution of the spectrum and our best estimate of the stellar velocity dispersion, together with the random and systematic errors. The mean $\bar{\sigma}$, the systematic uncertainty $\Delta\bar{\sigma}$, and the statistical uncertainty $\delta\bar{\sigma}$ of the velocity dispersion of an individual galaxy are defined as in KM25 such that
@@
where the index $k$ runs through the template library, $\delta\sigma_{k}$ is the statistical uncertainty estimated by pPXF, and $w_{k}$ is the weight we assigned.
Following KM25, we provide the stellar velocity dispersions and uncertainties as both the BIC-weighted average and equally weighted average of the measurements from two template libraries. The BIC-determined weights were estimated at the sample level to minimize noise in the process, and the samples were grouped by instrument and S/N. The estimated weights for the groups are summarized in Table 1. The preference for one of the libraries changed depending on the instrument and the S/N. For each galaxy, we used the weight for the group to which it was affiliated to measure the BIC-weighted average values. For example, when a galaxy was observed with LRIS and had an S/N>>20/Å, the weights we assigned to the measurements from Indo-US and MILES were 0.78 and 0.22, respectively. In general, the BIC-weighted and equally weighted velocity dispersions are very similar, which is consistent with the good agreement between the template libraries. In Fig. 7 we show the BIC-weighted velocity dispersions with the associated total uncertainties of the SL2S sample as a function of redshift.
When the number of template libraries is small (as in the case considered here), it is prudent to adopt Bessel’s correction to obtain the unbiased estimators (see KM25 for a derivation). Thus, the systematic uncertainties were Bessel corrected following
@@
For $k=2$, however, the Bessel correction nullifies the dependence on the weights, and $\Delta_{B}\bar{\sigma}$ reduces to $(\Delta_{B}\bar{\sigma})^{2}=(\frac{\sigma_{1}-\sigma_{2}}{2})^{2}$, where $\sigma_{1}$, and $\sigma_{2}$ are measurements using two template libraries. Thus, the Bessel-corrected systematic uncertainties, $\Delta_{B}\bar{\sigma}$, for the BIC-weighted and equally weighted velocity dispersions are equal and are reported as a single column in Table LABEL:table:veldis_measurements.
We also measured the BIC-weighted average statistical and systematic uncertainty, both Bessel corrected and uncorrected, for subsamples grouped by S/N and/or instrumental setup. In addition, we estimated the covariance matrix as follows:
@@
where the indices $i,j$ identify the elements of the covariance matrix, $\bar{\sigma}_{i,j}$ is the weighted-average velocity dispersion of the element at index $i$ or $j$, and $\delta_{i,j}$ is the Kronecker delta function. As above, the index $k$ runs through the template library, $\delta\bar{\sigma}_{i}$ is the weighted-average statistical uncertainty, and $w_{k}$ is the weight we assigned. Along the diagonal, the first term is the systematic error, and the second term is the statistical error. With Bessel correction, the estimator reads
@@
The normalized sample average uncertainties and the average amplitude of the off-diagonal terms of the covariance matrix between elements are summarized in Table 1. As expected, the random error is driven by the S/N, regardless of the instrumental setup, and it ranges between 3-8%. The systematic errors also depend on the S/N and instrumental setup, and they lie at a level of 1-2% without Bessel correction and at about 2-3% with the correction. This is only the error associated with template choice. Other potential sources of error, such as the wavelength range and polynomial order, were ignored because they are negligible with respect to the contribution from templates.
The average off-diagonal amplitude of the covariance matrix for different subsamples is positive. This is expected as the templates yielding higher (lower) velocity dispersion do so consistently for all galaxies. This quantity ranges around 1-2% within the same instrumental setup, and it lies at the subpercent level for the different instrumental setups. The LRIS subsamples have slightly higher covariances than the combined Xshooter and DEIMOS samples.
We found that Indo-US library is preferred by the LRIS data, while MILES is preferred by the other. This emphasizes the importance of using several stellar libraries and calculating BIC-weighted averages. In this case, however, similar results were obtained when the templates were given equal weight, so that the difference in BIC weight had less impact.
As a consistency test of our error estimates, we compared the stellar velocity dispersion obtained for 11 systems that have multiple spectra with a sufficiently high resolution. The fiducial spectrum was the spectrum with the highest S/N. The other spectra may have significantly lower S/N and may cover wavelength ranges differing by 500-1000 Å from their fiducial counterpart. The difference in the S/N quality of the spectra is reflected in the random uncertainties listed in Table LABEL:table:veldis_measurements. The comparison is shown in Fig. 8, where the error bars represent the sum in quadrature of the random and systematic errors. The sample average of the ratio is $\frac{\sigma_{\mathrm{high\ S/N}}}{\sigma_{\mathrm{low\ S/N}}}-1>=-0.11\pm 2.70$ %. Repeated measurements with higher S/Ns are needed to carry out an even more stringent test of the systematics for different instrumental setups.